OpenStudy (idealist10):

Find the curvature of y=x^3+1.

OpenStudy (idealist10):

@satellite73 @thomaster @ganeshie8 @iambatman @mathstudent55 @freckles @perl @e.mccormick

OpenStudy (adamaero):

http://tutorial.math.lamar.edu/Classes/CalcIII/Curvature.aspx Ooughh, Calc III

OpenStudy (idealist10):

\[k(x)=\frac{ \left| y'' \right| }{ [1+(y')^2]^{3/2} }\]

OpenStudy (idealist10):

I know that y'=3x^2 and y"=6x. So I got\[k(x)=\frac{ 6x }{ (1+9x^4)^{3/2} }\]

OpenStudy (idealist10):

But the answer in the book is\[\frac{ 6\left| x \right| }{ (1+9x^4)^{3/2} }\]

OpenStudy (idealist10):

I don't know which answer is right. @ganeshie8 @thomaster

OpenStudy (e.mccormick):

How did you get your answer to be out of the | |?

OpenStudy (idealist10):

So the book's answer is right?

OpenStudy (e.mccormick):

Well, look a the formula you posted. There are tules for dealing with | |. They don't change for calculus...

OpenStudy (e.mccormick):

rules that is...

OpenStudy (idealist10):

So the book's answer must be right then, is that correct?

OpenStudy (e.mccormick):

If that formula is correct, yes. Because you can't just take a variable outside of absolute value.

OpenStudy (idealist10):

Okay, thank you!